# AniMer+: Unified Pose and Shape Estimation Across Mammalia and Aves via Family-Aware Transformer

### v1

@@ III-B The Architecture of AniMer
Given an image $I\in\mathbb{R}^{H\times W\times 3}$ and a class token $\mathbf{x}_{\text{class}}\in\mathbb{R}^{1\times 1280}$, we first utilize a ViT encoder to extract image feature tokens $\mathbf{F}\in\mathbb{R}^{192\times 1280}$, and the class token interacts with the image to capture information about the animal family. We then feed the feature tokens $\mathbf{F}$ into a SMAL Transformer decoder to obtain a feature vector $\bm{f}\in\mathbb{R}^{1\times 1024}$. Finally, independent multi-layer perceptrons (MLPs) are used to predict the shape parameters $\hat{\beta}$, pose parameters $\hat{\theta}$, and camera parameters $\hat{\pi}$, where $\hat{\cdot}$ means estimated parameters. Meanwhile, the class token is fed into the predictor head for animal family supervised contrastive learning, which will be detailed in Sec. III-C. Note that we borrow the pretrained weights of the ViT encoder from ViTPose++ [26], which significantly enhances mesh recovery performance.
Despite the class token, AniMer features two key differences compared with previous HMR2.0 and HaMeR [16]. First, both HMR2.0 and HaMeR regress the residual SMPL/MANO parameters with respect to the non-zero mean parameters computed from large-scale motion databases. In contrast, AniMer directly decodes the final SMAL parameters due to limited SMAL pose prior. Second, both HMR2.0 and HaMeR train on the whole datasets in a single stage. Instead, we train AniMer in two stages. At the first stage, we train AniMer using only 3D data to ensure the network can predict feasible shapes and poses. At the second stage, all 3D and 2D data are introduced for training. The insight is that the size of 3D animal datasets is much smaller than that of 2D animal datasets, resulting in an imbalanced 3D-versus-2D data ratio.
@@ III-C Animal Family Supervised Contrastive Learning
Different from human whose SMPL shape parameters come from the same multivariate normal distribution, animals demonstrate at least two levels of shape differences: inter-family level and intra-family level. For example, dogs share similar shapes between each other, yet have distinct shapes from cows. To capture such two levels of shape distributions in AniMer, we propose an animal family loss based on supervised contrastive learning [61].
Specifically, we employ a learnable class token [48] to represent the animal family. First, this token $\mathbf{x}_{\text{class}}$ together with a minibatch of images $I\in\mathbb{R}^{B\times 3\times H\times W}$ are fed into the ViT encoder, where $B$ is the batch size. Then, an MLP head is utilized to generate the animal family feature $\bm{z}$ from the class token. Finally, we directly apply supervised contrastive learning to $\bm{z}$ through a loss $\mathcal{L}_{\text{con}}$ defined as follows:
@@
[…]
@@ III-D The Architecture of AniMer+
To address this challenge and create a truly unified model, we extend AniMer to AniMer+, which features a specialized architecture designed to concurrently handle multiple distinct taxa. The full architecture of AniMer+ is shown in Fig. 2. We utilize the idea of MoE [24, 25, 26]. Specifically, we partition the second FC layer into two parallel paths: two taxa-specific layers and one taxa-shared layer. The taxa-specific layers act as experts, learning to extract features relevant only to mammals or birds, respectively, while the shared layer processes features from all taxa, which incentivizes it during training to learn generalizable, cross-species representations.
@@
where $\mathbf{F}_{\text{taxa-shared}}\in\mathbb{R}^{193\times 960}$, $\mathbf{F}_{\text{taxa-specific}}\in\mathbb{R}^{193\times 320}$ and $\mathbf{F}_{\text{block}}\in\mathbb{R}^{193\times 1280}$. After the final ViT-MoE block, the feature tokens are passed to the decoder, which then branches to predict the distinct parameters for the SMAL (mammalia) and AVES (aves) models, respectively.
In this way, our method can effectively accommodate both mammals and birds. However, the existing bird datasets (e.g., CUB [62]) only provide 2D annotations, and the absence of 3D priors in single-view reconstruction tasks leads to depth ambiguities. In such context, we utilize our dataset generation pipeline (Sec. IV) to obtain the first bird dataset with 3D annotations named CtrlAVES3D. By combining AniMer+ with the dataset generation pipeline, our method can effectively accommodate both mammals and birds, with the potential to generalize to any animal group representable by a parametric model.
@@ III-E Loss Functions
$\mathcal{L}_{\text{3D}}=$ $\lambda_{\beta}||\hat{\beta}-\beta||_{2}^{2}+\lambda_{\theta}||\hat{\theta}-\theta||_{2}^{2}$ (4)
$+$ $||\hat{K}_{3D}-K_{3D}||_{1}+\lambda_{\alpha}||\hat{\alpha}-\alpha||_{1},$
@@
where $\lambda_{\beta}=0.5$, $\mu_{\beta}$, $\Sigma_{\beta}$, $\mu_{\theta}$, and $\Sigma_{\theta}$ are the mean and covariance of the prior distributions given by SMAL [1].
AVES Prior Loss. Following the AVES model [2], the prior loss for aves is as follows:
$\mathcal{L}_{\text{aves\_prior}}=\lambda_{\beta}||\beta||_{2}^{2}+\lambda_{\theta}||\hat{\theta}-\overline{\theta}||_{2}^{2}+||\hat{\alpha}-\overline{\alpha}||_{2}^{2},$ (7)
@@ IV Dataset Generation
SMAL Structure Condition for CtrlAni3D. The generation pipeline for the CtrlAni3D dataset is illustrated in Fig. 3. Given a posed SMAL mesh and viewpoint, we render it into a mask map and a depth map as the condition images to guide the structure of the images generated by a pre-trained ControlNet [22]. To guarantee the diversity of the poses and shapes, we randomly sample $\beta$ from the Gaussian distributed shape space provided by SMAL [1], and sample a more diverse range of $\theta$ from a combined pose space presented by dog [8, 11] and horse motion [63]. This is reasonable because the quadrupeds expressed by SMAL share similar anatomical structures. About the viewpoint for rendering, each dimension of the global rotation vector is uniformly sampled from $(-\pi,\pi)$ while the position is uniformly sampled between [0.5,0.5,4][-0.5,-0.5,4][ - 0.5 , - 0.5 , 4 ] and [0.5,0.5,8][0.5,0.5,8][ 0.5 , 0.5 , 8 ], therefore truncated images may be generated to enhance training’s robustness.
AVES Structure Condition for CtrlAVES3D. The primary differences in generating CtrlAVES3D compared to CtrlAni3D are three folds. 1) Instead of the early version ControlNet utilized by CtrlAni3D, we employ Flux-Union-Pro-V2 [64], a modern version ControlNet that generates more realistic bird textures. 2) Except for the shape parameter $\beta$ and the pose parameter $\theta$, the AVES model needs an extra bone parameter $\alpha$ to control the bone lengths. We sample $\alpha$ within [-0.5, 3.5] according to AVES’s practice [2]. 3) We utilize Canny edge and depth map as structural conditions, as Flux-Union-Pro-V2 does not support mask as a condition.
Text Condition. To further control the style of the generated images $x$, we seek to use text prompts. During the generation of the animal datasets of interest, we prompt ControlNet using common names instead of scientific names. For the SMAL parametric model, we manually classify the sampled 3D meshes into 10 species: cat, tiger, lion, cheetah, dog, wolf, horse, zebra, cow, and hippo, and use the species name as one keyword. For the AVES parametric model, the species name can be found in Table II. When it is hard to distinguish (e.g., some complicated cases of zebra/horse), we use both animal names as prompts and select the one with the highest cycle consistency. The second keyword is the pose description, e.g.,“stands” in Fig. 3, which is manually assigned according to the 3D mesh pose. Based on these keywords, ChatGPT [65] is employed to automatically complete a prompt sentence $c_{t}$ depicting possible animal behavior. Finally, both $c_{v}$ and $c_{t}$ act as the prompts of ControlNet for realistic and rich image generation.
@@
Semi-Automated Filtering. Note that not all generated images align perfectly with the conditions. To address this issue, we develop a semi-automated filtering strategy to reduce the burden on annotators. First, we utilize SAM2 [23] to extract the foreground mask from each generated image, enabling cycle-consistency checking by comparing it to the mask. Further, we manually filter out images that do not match the mesh poses to ensure high data quality. These misaligned results are denoted as “Failure cases”. Some failure and successful cases are illustrated in Fig. 4. Each synthetic image includes well-aligned annotations for the parametric model parameters ($\beta$, $\theta$, and $\gamma$ for SMAL; $\beta$, $\theta$, $\alpha$, and $\gamma$ for AVES) and 3D keypoints (26 keypoints for SMAL; 18 keypoints for AVES). By comparing the rendered depth image with the projected keypoint depths, we also obtain visible 2D keypoints as annotations. The visibility of the 2D keypoints is determined by comparing the depth $d_{k}$ of each keypoint with the depth $d_{p}$ at the corresponding pixel location. Specifically, visibility is assigned a value of 1 when $d_{k}\leq d_{p}$; otherwise, it is assigned a value of 0.
@@
Backgrounds. To ensure background diversity, we randomly sample COCO [20] images to serve as backgrounds for 2/72/72 / 7 of the CtrlAni3D images and 1/41/41 / 4 of the CtrlAVES3D images. The COCO backgrounds are used only when SAM2 achieves satisfactory segmentation quality. Therefore, increasing the ratio of the COCO backgrounds is equivalent to lowering the IoU threshold, which results in a decline in data quality. This may negatively impact model training.
@@ V-A Setup
Mammalian Datasets Summary. We curate and aggregate multiple mammalian datasets containing 2D and 3D annotations The full training dataset includes the training split of Animal Pose [50], APT-36K [66], AwA-Pose [67], Stanford Extra [11], Zebra synthetic [6], Animal3D [18], and our own CtrlAni3D. Only Animal3D, CtrlAni3D and Zebra synthetic have 3D annotations. For all datasets, we exclude images of animals not included in SMAL [1], such as elephants. Furthermore, we map the keypoints’ order across different datasets to be the same as that of the Animal3D dataset because the order of the annotated keypoints varies from dataset to dataset. For keypoints that are labeled in the Animal3D dataset but not labeled in a specific dataset, we set the weight of the keypoint to zero. We then aggregate all the aforementioned datasets (excluding Animal Kingdom [27]) for training, assigning different sampling weights to each dataset based on its type and size, as shown in Table III. For evaluation, we mainly use the test part of Animal3D, CtrlAni3D, and Animal Kingdom. Note that, Animal Kingdom serves as the OOD benchmark because it contains very challenging in-the-wild images which are never seen during training. A total of 8 mammalian species (horse, wolf, cheetah, leopard, dog, lion, jaguarundi cat and snow leopard) from Animal Kingdom are selected for testing.
@@
Avian Datasets Summary. We use the the training subset of the CUB dataset [62] and the training part of our CtrlAVES3D dataset as the training dataset. Note that only CtrlAVES3D has 3D annotations. Other than the CUB and CtrlAVES3D datasets, we also evaluate our model on the Cow Bird dataset [28] to demonstrate the generalizability of our model.
@@
Animal Pose Dataset. The Animal Pose dataset [50] includes 5 categories: dog, cat, cow, horse and sheep, comprising a total of over 6,000 instances across more than 4,000 images. Each animal instance is annotated with 20 keypoints.
@@
APT-36k Dataset. The APT-36k dataset [66] contains 36000 images covering 30 animal species from different scenes. Each animal instance is labeled with 17 keypoints.
@@
AwA Pose Dataset. The AwA Pose dataset [67] is introduced for 2D quadruped animal pose estimation. It contains 10064 images of 35 quadruped animal species and each image is annotated with 39 keypoints.
@@
Stanford Extra Dataset. The Stanford Extra dataset [11] consists of 20,580 images and covers 120 dog breeds. Each image is annotated with 20 2D keypoints and silhouette.
@@
Zebra synthetic Dataset. The Zebra synthetic dataset [6] consists of 12850 images. Each image is randomly generated that differs in background, shape, pose, camera, and appearance.
@@
Animal Kingdom Dataset. The Animal Kingdom dataset [27] includes a diverse range of animal species. We only use 8 major mammalian species within the pose estimation dataset to evaluate our method.
@@
Animal3D Dataset. The Animal3D dataset [18] contains a total of 3379 images, which are classified into 40 classes. Each image is annotated with SMAL [1] parameters, 2D keypoints, 3D keypoints and masks.
@@
CUB Dataset. The CUB dataset [62] contains 200 subcategories of birds. Each image is labeled with a subcategory label, 15 part locations, 312 binary attributes and a bounding box.
CUB17 Dataset. The CUB17 dataset [2] contains 17 categories selected from the CUB200 dataset, details of which are presented in Table II. Different from the CUB dataset, images in the CUB17 dataset are annotated with 18 keypoints.
@@
Cow Bird Dataset. The Cow Bird dataset [28] comprises 1,000 images of 15 brown-headed cowbirds recorded at 40 Hz during a three-month mating season, annotated with 6,355 pixel-wise segmentation masks and corresponding bounding boxes. A subset of 18 moments includes 1,031 mask annotations, each enriched with 12 semantic keypoints per bird.
@@
AwA-Pose 2884 5.4% 0.15
Zebra Synthetic 12850 23.9% 0.05
Stanford Extra 7689 14.3% 0.15
APT-36K 4887 9.1% 0.15
@@
Training Details. Our model is implemented by PyTorch [68]. We use AdamW [69] optimizer with a linear learning rate decay schedule. The initial learning rate is $1.25\times 10^{-6}$. For the mammalian experiment, we train AniMer-a (replacing the ViT with ResNet152), AniMer-b (no pretraining), AniMer-c (replacing the Transformer decoder with MLP), and AniMer-M (training AniMer on mammalian datasets) for 2M steps. For the avian experiments, we train AniMer-A (training AniMer on avian datasets) for 240K steps. To directly compare AniMer and AniMer+, we train AniMer-AM (training AniMer on both mammalian and avian datasets) and AniMer+ for 1.4M steps on the same datasets. All experiments above are conducted on a single NVIDIA RTX 4090 GPU.
@@
TABLE V: Quantitative comparisons on the Animal3D dataset with 3D Fauna [37]. PA-CD: Chamfer Distance after Procrustes alignment between the 3D grouth truth and predicted results. Results for PCK@0.1 of 3D Fauna are borrowed from the original paper.

### v2

@@ III-B The Architecture of AniMer
Given an image $I\in\mathbb{R}^{H\times W\times 3}$ and a class token $\mathbf{x}_{\text{class}}\in\mathbb{R}^{1\times 1280}$, we first utilize a ViT encoder to extract image feature tokens $\mathbf{F}\in\mathbb{R}^{192\times 1280}$, and the class token interacts with the image to capture information about the animal family. We then feed the feature tokens $\mathbf{F}$ into a SMAL Transformer decoder to obtain a feature vector $\bm{f}\in\mathbb{R}^{1\times 1024}$. Finally, independent multi-layer perceptrons (MLPs) are used to predict the shape parameters $\hat{\beta}$, pose parameters $\hat{\theta}$, and camera parameters $\hat{\pi}$, where $\hat{\cdot}$ means estimated parameters. Meanwhile, the class token is fed into the predictor head for animal family supervised contrastive learning, which will be detailed in Sec. III-C. Note that we borrow the pretrained weights of the ViT encoder from ViTPose++ [xu2023vitpose++], which significantly enhances mesh recovery performance.
Despite the class token, AniMer features two key differences compared with previous HMR2.0 and HaMeR [pavlakos2024reconstructing]. First, both HMR2.0 and HaMeR regress the residual SMPL/MANO parameters with respect to the non-zero mean parameters computed from large-scale motion databases. In contrast, AniMer directly decodes the final SMAL parameters due to limited SMAL pose prior. Second, both HMR2.0 and HaMeR train on the whole datasets in a single stage. Instead, we train AniMer in two stages. At the first stage, we train AniMer using only 3D data to ensure the network can predict feasible shapes and poses. At the second stage, all 3D and 2D data are introduced for training. The insight is that the size of 3D animal datasets is much smaller than that of 2D animal datasets, resulting in an imbalanced 3D-versus-2D data ratio.
@@ III-C Animal Family Supervised Contrastive Learning
Different from human whose SMPL shape parameters come from the same multivariate normal distribution, animals demonstrate at least two levels of shape differences: inter-family level and intra-family level. For example, dogs share similar shapes between each other, yet have distinct shapes from cows. To capture such two levels of shape distributions in AniMer, we propose an animal family loss based on supervised contrastive learning [khosla2020supervised].
Specifically, we employ a learnable class token [dosovitskiy2020image] to represent the animal family. First, this token $\mathbf{x}_{\text{class}}$ together with a minibatch of images $I\in\mathbb{R}^{B\times 3\times H\times W}$ are fed into the ViT encoder, where $B$ is the batch size. Then, an MLP head is utilized to generate the animal family feature $\bm{z}$ from the class token. Finally, we directly apply supervised contrastive learning to $\bm{z}$ through a loss $\mathcal{L}_{\text{con}}$ defined as follows:
@@
Within a minibatch, $P(i)$ represents samples that share the same family label $i$, $O(i)$ represents samples other than $i$. The notation $|P(i)|$ denotes the size of this set. $\tau\in\mathbb{R}^{+}$ is a scalar temperature parameter.
@@ III-D The Architecture of AniMer+
To address this challenge and create a truly unified model, we extend AniMer to AniMer+, which features a specialized architecture designed to concurrently handle multiple distinct taxa. The full architecture of AniMer+ is shown in Fig. 2. We utilize the idea of MoE [shazeer2017outrageously, jiang2024mixtral, xu2023vitpose++]. Specifically, we partition the second FC layer into two parallel paths: two taxa-specific layers and one taxa-shared layer. The taxa-specific layers act as experts, learning to extract features relevant only to mammals or birds, respectively, while the shared layer processes features from all taxa, which incentivizes it during training to learn generalizable, cross-species representations.
@@
where $\mathbf{F}_{\text{taxa-shared}}\in\mathbb{R}^{193\times 960}$, $\mathbf{F}_{\text{taxa-specific}}\in\mathbb{R}^{193\times 320}$ and $\mathbf{F}_{\text{block}}\in\mathbb{R}^{193\times 1280}$. After the final ViT-MoE block, the feature tokens are passed to the decoder, which then branches to predict the distinct parameters for the SMAL (mammalia) and AVES (aves) models, respectively. Different from the typical MoE where a gating network is necessary to predict which expert to route each token to, we use a simple hard-coded switch based on the input’s taxa (mammals or birds), which is more direct and efficient than a learned gate for this problem.
In this way, our method can effectively accommodate both mammals and birds. However, the existing bird datasets (e.g., CUB [wah2011caltech]) only provide 2D annotations, and the absence of 3D priors in single-view reconstruction tasks leads to depth ambiguities. In such context, we utilize our dataset generation pipeline (Sec. IV) to obtain the first bird dataset with 3D annotations named CtrlAVES3D. By combining AniMer+ with the dataset generation pipeline, our method can effectively accommodate both mammals and birds, with the potential to generalize to any animal group representable by a parametric model.
@@ III-E Loss Functions
[…]
@@
where $\lambda_{\beta}=0.5$, $\mu_{\beta}$, $\Sigma_{\beta}$, $\mu_{\theta}$, and $\Sigma_{\theta}$ are the mean and covariance of the prior distributions given by SMAL [zuffi20173d].
AVES Prior Loss. Following the AVES model [wang2021birds], the prior loss for aves is as follows:
[…]
@@ IV Dataset Generation
SMAL Structure Condition for CtrlAni3D. The generation pipeline for the CtrlAni3D dataset is illustrated in Fig. 3. Given a posed SMAL mesh and viewpoint, we render it into a mask map and a depth map as the condition images to guide the structure of the images generated by a pre-trained ControlNet [zhang2023adding]. To guarantee the diversity of the poses and shapes, we randomly sample $\beta$ from the Gaussian distributed shape space provided by SMAL [zuffi20173d], and sample a more diverse range of $\theta$ from a combined pose space presented by dog [ruegg2023bite, biggs2020wldo] and horse motion [li2024poses]. This is reasonable because the quadrupeds expressed by SMAL share similar anatomical structures. About the viewpoint for rendering, each dimension of the global rotation vector is uniformly sampled from $(-\pi,\pi)$ while the position is uniformly sampled between [0.5,0.5,4][-0.5,-0.5,4] and $[0.5,0.5,8]$, therefore truncated images may be generated to enhance training’s robustness.
AVES Structure Condition for CtrlAVES3D. The primary differences in generating CtrlAVES3D compared to CtrlAni3D are four folds. 1) Instead of the early version ControlNet utilized by CtrlAni3D, we employ Flux-Union-Pro-V2 [flux-cn-union-pro-2], a modern version ControlNet that generates more realistic bird textures. 2) Due to the lack of a probabilistic distribution of pose or shape space, we randomly sample $\beta$ and $\theta$ from the well-estimated results provided by the original AVES paper [wang2021birds] on the 297 images of the CUB17 dataset (see Sec. V-A). 3) Except for the shape parameter $\beta$ and the pose parameter $\theta$, the AVES model needs an extra bone parameter $\alpha$ to control the bone lengths. We sample $\alpha$ within [-0.5, 3.5] according to AVES’s practice [wang2021birds]. 4) We utilize Canny edge and depth map as structural conditions, as Flux-Union-Pro-V2 does not support mask as a condition.
Text Condition. To further control the style of the generated images $x$, we seek to use text prompts. During the generation of the animal datasets of interest, we prompt ControlNet using common names instead of scientific names. For the SMAL parametric model, we manually classify the sampled 3D meshes into 10 species: cat, tiger, lion, cheetah, dog, wolf, horse, zebra, cow, and hippo, and use the species name as one keyword. For the AVES parametric model, the species name can be found in Table II. When it is hard to distinguish (e.g., some complicated cases of zebra/horse), we use both animal names as prompts and select the one with the highest cycle consistency. The second keyword is the pose description, e.g.,“stands” in Fig. 3, which is manually assigned according to the 3D mesh pose. Based on these keywords, ChatGPT [achiam2023gpt] is employed to automatically complete a prompt sentence $c_{t}$ depicting possible animal behavior. Finally, both $c_{v}$ and $c_{t}$ act as the prompts of ControlNet for realistic and rich image generation.
@@
Semi-Automated Filtering. Note that not all generated images align perfectly with the conditions. To address this issue, we develop a semi-automated filtering strategy to reduce the burden on annotators. First, we utilize SAM2 [ravi2024sam2] to extract the foreground mask from each generated image, enabling cycle-consistency checking by comparing it to the mask. Further, we manually filter out images that do not match the mesh poses to ensure high data quality. These misaligned results are denoted as “Failure cases”. Some failure and successful cases are illustrated in Fig. 4. Each synthetic image includes well-aligned annotations for the parametric model parameters ($\beta$, $\theta$, and $\gamma$ for SMAL; $\beta$, $\theta$, $\alpha$, and $\gamma$ for AVES) and 3D keypoints (26 keypoints for SMAL; 18 keypoints for AVES). By comparing the rendered depth image with the projected keypoint depths, we also obtain visible 2D keypoints as annotations. These are feasible because all the keypoints are defined on the body surface. Specifically, the visibility of the 2D keypoints is determined by comparing the depth $d_{k}$ of each keypoint with the depth $d_{p}$ at the corresponding pixel location. Specifically, visibility is assigned a value of 1 when $d_{k}\leq d_{p}$; otherwise, it is assigned a value of 0.
@@
Backgrounds. To ensure background diversity, we randomly sample COCO [lin2014microsoft] images to serve as backgrounds for $2/7$ of the CtrlAni3D images and $1/4$ of the CtrlAVES3D images. The COCO backgrounds are used only when SAM2 achieves satisfactory segmentation quality. Therefore, increasing the ratio of the COCO backgrounds is equivalent to lowering the IoU threshold, which results in a decline in data quality. This may negatively impact model training.
@@ V-A Setup
Mammalian Datasets Summary. We curate and aggregate multiple mammalian datasets containing 2D and 3D annotations The full training dataset includes the training split of Animal Pose [cao2019cross], APT-36K [yang2022apt], AwA-Pose [banik2021novel], Stanford Extra [biggs2020wldo], Zebra synthetic [zuffi2019three], Animal3D [xu2023animal3d], and our own CtrlAni3D. Only Animal3D, CtrlAni3D and Zebra synthetic have 3D annotations. For all datasets, we exclude images of animals not included in SMAL [zuffi20173d], such as elephants. Furthermore, we map the keypoints’ order across different datasets to be the same as that of the Animal3D dataset because the order of the annotated keypoints varies from dataset to dataset. For keypoints that are labeled in the Animal3D dataset but not labeled in a specific dataset, we set the weight of the keypoint to zero. We then aggregate all the aforementioned datasets (excluding Animal Kingdom [Ng_2022_CVPR]) for training, assigning different sampling weights to each dataset based on its type and size, as shown in Table III. For evaluation, we mainly use the test part of Animal3D, CtrlAni3D, and Animal Kingdom. Note that, Animal Kingdom serves as the OOD benchmark because it contains very challenging in-the-wild images which are never seen during training. A total of 8 mammalian species (horse, wolf, cheetah, leopard, dog, lion, jaguarundi cat and snow leopard) from Animal Kingdom are selected for testing.
@@
Avian Datasets Summary. We use the the training subset of the CUB dataset [wah2011caltech] and the training part of our CtrlAVES3D dataset as the training dataset. Note that only CtrlAVES3D has 3D annotations. Other than the CUB and CtrlAVES3D datasets, we also evaluate our model on the Cow Bird dataset [badger20203d] to demonstrate the generalizability of our model.
@@
Animal Pose Dataset. The Animal Pose dataset [cao2019cross] includes 5 categories: dog, cat, cow, horse and sheep, comprising a total of over 6,000 instances across more than 4,000 images. Each animal instance is annotated with 20 keypoints.
@@
APT-36k Dataset. The APT-36k dataset [yang2022apt] contains 36000 images covering 30 animal species from different scenes. Each animal instance is labeled with 17 keypoints.
@@
AwA Pose Dataset. The AwA Pose dataset [banik2021novel] is introduced for 2D quadruped animal pose estimation. It contains 10064 images of 35 quadruped animal species and each image is annotated with 39 keypoints.
@@
Stanford Extra Dataset. The Stanford Extra dataset [biggs2020wldo] consists of 20,580 images and covers 120 dog breeds. Each image is annotated with 20 2D keypoints and silhouette.
@@
Zebra synthetic Dataset. The Zebra synthetic dataset [zuffi2019three] consists of 12850 images. Each image is randomly generated that differs in background, shape, pose, camera, and appearance.
@@
Animal Kingdom Dataset. The Animal Kingdom dataset [Ng_2022_CVPR] includes a diverse range of animal species. We only use 8 major mammalian species within the pose estimation dataset to evaluate our method.
@@
Animal3D Dataset. The Animal3D dataset [xu2023animal3d] contains a total of 3379 images, which are classified into 40 classes. Each image is annotated with SMAL [zuffi20173d] parameters, 2D keypoints, 3D keypoints and masks.
@@
CUB Dataset. The CUB dataset [wah2011caltech] contains 200 subcategories of birds. Each image is labeled with a subcategory label, 15 part locations, 312 binary attributes and a bounding box.
CUB17 Dataset. The CUB17 dataset [wang2021birds] contains 17 categories (as listed in Table II) selected from the CUB dataset and 297 images in total. Different from the CUB dataset, the images in the CUB17 dataset are annotated with 18 keypoints. This dataset is only used for training, and the AVES models estimated by Wang et al. [wang2021birds] on these images are used as the pose and shape space of CtrlAVES3D.
@@
Cow Bird Dataset. The Cow Bird dataset [badger20203d] comprises 1,000 images of 15 brown-headed cowbirds recorded at 40 Hz during a three-month mating season, annotated with 6,355 pixel-wise segmentation masks and corresponding bounding boxes. A subset of 18 moments includes 1,031 mask annotations, each enriched with 12 semantic keypoints per bird.
@@
AwA-Pose 2884 5.3% 0.15
Zebra Synthetic 12850 23.8% 0.05
Stanford Extra 7689 14.2% 0.15
APT-36K 4887 9.0% 0.15
@@
Training Details. Our model is implemented by PyTorch [paszke2019pytorch]. We use AdamW [adamw] optimizer with a linear learning rate decay schedule. The initial learning rate is $1.25\times 10^{-6}$. For mammalian experiments, we train AniMer-a (replacing the ViT with ResNet152), AniMer-b (no pretraining), AniMer-c (replacing the Transformer decoder with MLP), and AniMer-M (training AniMer on mammalian datasets) for total 2M steps. Specifically, at the first stage, these models are trained for 200K steps (500 epochs) using only 3D datasets to establish a strong 3D prior. At the second stage, these models are then fine-tuned for an additional 1.8M steps (700 epochs) on the aggregated collection of all 3D and 2D datasets.
For avian experiments, we train AniMer-A (training AniMer on avian datasets) for 240K steps in a single stage.
To directly compare AniMer and AniMer+, we train AniMer-AM (training AniMer on both mammalian and avian datasets) and AniMer+ for 1.4M steps on the same datasets. Except for specifically stated experiments, all other experimental setups are conducted under the aforementioned conditions. All experiments above are conducted on a single NVIDIA RTX 4090 GPU.
@@
TABLE V: Quantitative comparisons on the Animal3D dataset with 3D Fauna [li2024learning]. PA-CD: Chamfer Distance after Procrustes alignment between the 3D grouth truth and predicted results. Results for PCK@0.1 of 3D Fauna are borrowed from the original paper.